OpenStudy (anonymous):

Help please! Medal for explanation! Use the arc length formula and the given information to find r. Show your work for full credit. s = 8.1 ft θ = π/3 rad r = ?

OpenStudy (anonymous):

@myininaya

OpenStudy (anonymous):

This is pre-calculus...

OpenStudy (anonymous):

s = 8.1 ft θ = π/3 rad r =

OpenStudy (philcua):

Ahh nice bunny

OpenStudy (philcua):

Not making a joke actually nice bunny

OpenStudy (anonymous):

@hartnn can u please help me?

hartnn (hartnn):

sure you know the formula for arc length ?

OpenStudy (anonymous):

No...

hartnn (hartnn):

ok, here it is, quite simple \(s = r \theta \) you know s and theta just plug them in!

hartnn (hartnn):

\(8.1 = r (\pi/3) \) find r :)

OpenStudy (anonymous):

Okay, one sec...

OpenStudy (anonymous):

8.1/θ = r ?

hartnn (hartnn):

yes plug in theta and find r

OpenStudy (anonymous):